# A Call for New Recipes to Enhance Spatial Reasoning in MLLMs

### v1

@@ 3.3 Training Objective
In this section, we explore the training dynamics of MLLMs, emphasizing their current limitations in addressing spatial reasoning. We begin by introducing the loss function design and outlining the training modalities where the loss is computed. Subsequently, we highlight the shortcomings of the existing training paradigms in improving spatial reasoning.
@@
MLLMs with Discrete Visual Tokens. MLLMs with discrete visual tokens use visual tokenizer to project visual features into discrete token representations. Visual tokens are then fed into language model backbone along with language tokens for reasoning. However, this integration to Transformer backbone primarily focuses on semantic reasoning, which is optimized by cross-entropy loss, without explicitly addressing the nuances of spatial reasoning. This approach presents several challenges as follows. 1) Mismatch in Token Distributions: Visual tokens are treated identically to text tokens. Unlike text tokens, which follow Zipf’s law (Piantadosi, 2014), visual tokens originate from an entirely different distribution and thereby are not yet to be determined to be inherently suitable for unified modeling with language tokens in the same way (Chan et al., 2024). 2) Absence of Spatial Alignment: Cross-entropy loss primarily aligns semantic content between text and visual tokens but fails to incorporate explicit signals for spatial layout understanding. Recent works, such as XVLM (Zeng et al., 2022), address this limitation by introducing localized captions and bounding boxes to facilitate fine-grained spatial-semantic alignment. However, these methods heavily rely on the granularity of annotated datasets, which may be influenced in cases where annotations are sparse or noisy. Premsri & Kordjamshidi (2024) have explored neuro-symbolic training methods that embed symbolic rules into loss design, aiming to mitigate the lack of spatial constraints. However, it remains constrained to the textual domain, lacking the multimodal context required for richer spatial understanding.
@@
MLLMs with Continuous Denoising Process. Continuous modeling, as seen in diffusion denoising approaches, introduces another novel paradigm of MLLMs for handling spatial reasoning. Notable advancements in this area include Transfusion (Zhou et al., 2024a) and LatentLM(Sun et al., 2024b), both of which combine discrete token modeling for text with a continuous diffusion process for generating images. While diffusion loss guides the denoising process, its effectiveness in capturing spatial layouts remains unclear without spatial constraints (e.g. bounding boxes) (Patel & Serkh, 2024). This hybrid approach highlights the potential of leveraging both discrete and continuous representations, but further studies are required to determine how well image-level representations retain detailed spatial relationship beyond the semantics of individual entities. Exploring spatially aware loss functions or additional supervisory signals in this context could provide significant insights.
@@
Training Modality. Another key aspect for training MLLMs lies in the modality where the loss is calculated. Most MLLMs calculate loss solely based on textual tokens, limiting the flow of information from other modalities. In contrast, recent advancements in multimodal generation incorporate image data into the optimization process, enabling richer cross-modal interactions. Focusing on dynamic spatial reasoning, Li et al. (2025) compare a text-only setup with an interleaved text-image scenario where images do not contribute to the loss. Interestingly, even when the model is optimized solely on text tokens, interleaved text-image training outperforms or matches text-only training. When supervision signals are derived from both text and visual tokens, the model demonstrates superior performance, generating more effective visual and verbal reasoning traces. This highlights the importance of leveraging multimodal information to enhance supervision signals, in addition to careful loss function design.
@@ 4 Evaluation Benchmark
Robust evaluation benchmarks are essential for advancing spatial reasoning capabilities in MLLMs, as they not only measure current abilities but also highlight critical gaps and guide future research directions. Several benchmarks (Mirzaee et al., 2021; Shi et al., 2022; Rizvi et al., 2024; Li et al., 2024b) focused on text-only scenarios, where models were evaluated purely through language descriptions. While these benchmarks provided initial insights, the emergence of MLLMs has necessitated more comprehensive evaluation frameworks that can assess spatial reasoning across different modalities and scenarios.
@@ 4.1 Static Spatial Reasoning
VSR (Liu et al., 2023) pioneered comprehensive evaluation across 66 spatial relation types, revealing significant gaps between early VLMs and human performance, and highlighting the crucial role of positional encodings. What’sUp (Kamath et al., 2023) in contrast, narrowed its scope to household object pairs in spatial relationships. It uncovered models’ preference for color attributes over spatial relationships, highlighting specific limitations in spatial understanding.
Beyond 2D tasks, SpatialVLM (Chen et al., 2024a) introduced 3D spatial VQA tasks incorporating both qualitative and quantitative assessments, while SpatialRGBT-Bench (Cheng et al., 2024) further advanced the evaluation landscape by providing ground-truth 3D annotations across diverse environments. Similarly, CV-Bench (Tong et al., 2024) evaluated the fundamental 2D (Spatial Relationship and Object count) and 3D (Depth Order and Relative Distance) visual understanding of MLLMs. Collectively, these benchmarks have demonstrated that using data with corresponding spatial annotations for fine-tuning can improve the model’s performance on associated tasks.
In addition to general benchmarks, other works have focused on more specialized tasks, yielding deeper insights. For instance, Wang et al. introduced three novel VQA benchmarks (Spatial-Map, Maze-Nav, and Spatial-Grid), revealing following findings: extra visual input offers limited improvement over MLLMs’ LLM backbone, and MLLMs tend to rely less on visual information when sufficient textual context is available. TopViewRS (Li et al., 2024a) specifically examined spatial reasoning from top-view perspectives, revealing a substantial performance gap between MLLMs and humans. Additionally, COMFORT (Zhang et al., 2024b) systematically assessed Frames of Reference (FoR) using 3D images. Their findings showed that current MLLMs struggle with robustness and consistency in FoR tasks. Specifically, most models exhibit a preference for reflected coordinate transformations and egocentric relative FoR but face challenges in adapting flexibly to other FoRs.
@@ 4.2 Dynamic Spatial Reasoning
The recent work SAT (Ray et al., 2024) goes beyond static questions to more dynamic tasks, including Object Movement, Goal Aiming, Action Consequence, Egocentric Movement and Allocentric Perspective. The results demonstrated that stronger closed-source models (GPT4-o (Achiam et al., 2023) and Gemini-1.5-pro (Team et al., 2023)) and spatially-tuned models (Robopoint (Yuan et al., 2024) and Cambrian-1 (Tong et al., 2024)) still struggle on dynamic tasks despite performing well on static. In particular, the perspective tasks involving Transformation Reasoning capability remain challenging even for strong models like GPT4-o.
Another recent study VSI-Bench (Yang et al., 2024) evaluated dynamic spatial reasoning tasks through a video-based benchmark. One of their key observations is that the best-performing MLLM use global spatial context and reasoning to infer correctly, indicating MLLMs may develop an implicit world model. And the error analysis reveals that almost 50% of errors stems from egocentric-allocentric transformation, indicating a lack of Transformation Reasoning capability. Additionally, linguistic prompting techniques like CoT (Wei et al., 2022), although effective in language reasoning and general visual tasks, are harmful for spatial reasoning tasks. They also leveraged cognitive maps (Tolman, 1948) to explore how MLLMs remember spaces, finding it forms a series of local world models from the video, rather than a unified global model. In addition, SPACE (Ramakrishnan et al., 2024) evaluates whether MLLMs demonstrate spatial cognition and finds that frontier models perform near chance level on several classic tests of animal cognition.
Despite the comprehensive evaluation of MLLMs’ relational and transformation reasoning abilities across various benchmarks, there remains a gap in correlating these abilities with real-world applications, such as Embodied AI. While current benchmarks provide valuable insights, they lack empirical and theoretical investigations that demonstrate how improvements in spatial reasoning directly translate to enhanced performance in applications. We believe that future research should bridge this gap, providing a deeper understanding of how stronger spatial reasoning in MLLMs can lead to better outcomes in practical applications.
Our position:  Existing benchmarks lack a direct correlation with real-world applications like Embodied AI. We advocate for an evaluation framework that provides empirical investigations to establish a connection between enhanced spatial reasoning in MLLMs and improved performance on diverse applications.
@@ 5 Reasoning Method
Figure 2: The responses and attention visualization of Qwen-2VL on path reading and planning tasks. The attention maps provide insights into the model’s spatial reasoning ability, highlighting how it attends to visual elements during task execution.
@@
Direct reasoning involves the model directly generating the answer. Instead of directly outputting the answer, recent studies have shown that scaling test-time compute (Snell et al., 2024b) improves reasoning performance. This is achieved through methods such as search algorithms or strategies like CoT reasoning (Wei et al., 2022; Zhang et al., 2023). Within spatial reasoning, the effectiveness of CoT reasoning has also been extensively explored. For example, Li et al. (2024a) demonstrated its utility in eliciting spatial reasoning. Besides, Wu et al. (2024) enhanced performance in grid-based dynamic spatial reasoning tasks by representing reasoning traces in an ASCII-style format. However, in multimodal spatial reasoning, textual descriptions can be verbose and sometimes imprecise in describing the spatial relationships in the image. Li et al. (2025) introduces a novel reasoning paradigm, MVoT, that extends reasoning traces from textual to multimodal representations. This opens new research direction and opportunities for multimodal native models to tackle the problem of spatial reasoning, with more straightforward illustrations during reasoning.
In addition to utilizing the reasoning capability of a single model, recent advancement in multimodal agentic reasoning unlocks the potential of incorporating multiple models, each acting as an expert in specific tasks, to improve spatial reasoning performance. For example, Image-of-Thought prompting (Zhou et al., 2024b) introduces a spatial ruler to accurately perceive spatial relationships and address spatial positioning queries. Similarly, Cai et al. (2024) enable vision-language models (VLMs) to call an external API for depth estimation, enhancing the understanding of spatial information in input images. MM-ReAct (Yang et al., 2023) employs vision expert models within a ReAct-style pipeline to gather observations that support decision-making. Despite of these efforts, research on designing agentic workflows specifically for spatial reasoning remains limited. Future directions could explore the agentic workflow tailored for spatial reasoning capabilities outlined in Section 2, for example, either by structuring the spatial relations with scene graph (Zhong et al., 2021), or by imagining the spatial transformation with MVoT-alike systems.
@@ 6 Case Study
To further illustrate the limitations of current MLLM, we designed two representative tasks, as illustrated in the top section of Figure 2. In the Path Reading task, the model is provided with an image of a maze and asked to describe the path in it. The Planning task goes a step further by requiring the model to plan a path from a given start point to a specified goal in the maze. The detailed prompts can be found in Appendix A. In this case study, we employ Qwen2-VL (Wang et al., 2024b) and LLaVa 1.5 (Liu et al., 2024). The results of Qwen2-VL are shown in Figure 2 with LLaVa 1.5 (Liu et al., 2024) results in Appendix A.
The goal of the Path Reading task is to evaluate whether the model can accurately attend to the corresponding regions in the image. Although both models fail to generate a complete correct answer, our findings reveal distinct attention patterns between the two models. Qwen-2VL, compared to LLaVA-1.5, demonstrates a more accurate attention mechanism, as its correct responses consistently focus on the relevant regions of the image. This difference may stem from architectural enhancements in Qwen-2VL, such as the use of 2D-RoPE in the vision encoder and M-RoPE in the language model backbone. These observations provide additional evidence supporting our analysis of position embedding in Sec 3.2.
The Planning task requires more complex spatial reasoning capabilities. Our results indicate that both Qwen-2VL and LLaVA-1.5 fail to generate correct path, while responses of Qwen-2VL are much more reasonable. The attention maps demonstrates that Qwen-2VL primarily focuses on object-level semantic information, such as ‘yellow star’, rather than forming a higher-level understanding of the spatial layout. To further explore these limitations, we introduced extra visual cues, as illustrated in the bottom section of Figure 2. Additionally, we employed the CoT prompting technique to investigate the reasoning steps. The results reveal that Qwen-2VL generates a reasonable path while ignoring the black boundary. As shown in attention maps, Qwen-2VL also attends to the visual cues during its reasoning process. However, it remains challenging when reasoning about the accurate spatial relationship, like ‘right to the red triangle’.
@@
7 Alternative Views
We would like to describe and address other alternative views to foster further discussions. Despite the prior discussion calling for new recipes to enhance spatial reasoning in MLLMs, potential alternative position may suggest that current recipe can emerge spatial reasoning capability. Possible direction to address this position includes keeping scaling up training data (Kaplan et al., 2020) or test-time compute (Snell et al., 2024a) within the current paradigm.
Research shows that MLLMs can understand basic spatial relationships (Li et al., 2024a), despite not being specifically optimized for spatial reasoning during training. It is reasonable to conjecture that with more data being included during training, the model can implicitly obtain the spatial awareness. This also corroborates with recent discussion of whether video model can be seen as a pathway towards world model (Kang et al., 2024), where the ‘world’ is represented by the model learnt from massive real-world videos without specifically injecting the physical rules.
In addition, scaling test-time compute has gained intense attention in improving reasoning performance without tuning parameters. Motivated by improvements in reasoning over relative spatial relations with CoT-style prompting (Wu et al., 2024; Li et al., 2024a), further research can explore better strategies for scaling test-time compute to cultivate the potential of current MLLMs without the new recipe introduced before.
@@ 8 Conclusion
This position paper discusses a critical and timely gap of spatial reasoning capabilities in MLLMs. By systematically dissecting the challenges in spatial reasoning, we underscore the urgent need for targeted research and development. Our analysis reveals that the components of current recipes-from training data to reasoning mechanisms-limit the spatial reasoning capability of MLLMs , not only constraining their real-world applicability but also hindering the way to true intelligence. Therefore, we call on the community to prioritize spatial reasoning and develop new recipe for MLLMs to effectively enhance their spatial reasoning capabilities.

### v2

@@ 3.3 Training Objective
In this section, we explore the training dynamics of MLLMs, emphasizing their limitations in addressing spatial reasoning. We first introduce the loss function design and outline the training modalities. Then, we highlight the shortcomings of training paradigms in improving spatial reasoning.
@@
MLLMs with Discrete Visual Tokens. MLLMs with discrete visual tokens use visual tokenizer to project visual features into discrete token representations. Visual tokens are then fed into language model backbone along with language tokens for reasoning. However, this integration to Transformer backbone primarily focuses on semantic reasoning, which is optimized by cross-entropy loss, without explicitly addressing the nuances of spatial reasoning. This approach presents several challenges as follows. 1) Mismatch in Token Distributions: Visual tokens are treated identically to text tokens. Unlike text tokens, which follow Zipf’s law [78], visual tokens originate from an entirely different distribution and thereby are not yet to be determined to be inherently suitable for unified modeling with language tokens in the same way [79]. 2) Absence of Spatial Alignment: Cross-entropy loss primarily aligns semantic content between text and visual tokens but fails to incorporate explicit signals for spatial layout understanding. Recent works, such as XVLM [49], address this limitation by introducing localized captions and bounding boxes to facilitate fine-grained spatial-semantic alignment. However, these methods heavily rely on the granularity of annotated datasets, which may be influenced in cases where annotations are sparse or noisy. Premsri and Kordjamshidi [80] have explored neuro-symbolic training methods that embed symbolic rules into loss design, aiming to mitigate the lack of spatial constraints. However, it remains constrained to the textual domain, lacking the multimodal context required for richer spatial understanding.
@@
MLLMs with Continuous Denoising Process. Continuous modeling, as seen in diffusion denoising approaches, introduces another novel paradigm of MLLMs for handling spatial reasoning. Notable advancements in this area include Transfusion [81] and LatentLM[82], both of which combine discrete token modeling for text with a continuous diffusion process for generating images. While diffusion loss guides the denoising process, its effectiveness in capturing spatial layouts remains unclear without spatial constraints (e.g. bounding boxes) [83]. This hybrid approach highlights the potential of leveraging both discrete and continuous representations, but further studies are required to determine how well image-level representations retain detailed spatial relationship. Exploring spatially aware loss functions or additional supervisory signals could provide significant insights.
@@
Training Modality. Another key aspect for training MLLMs lies in the modality where the loss is calculated. Most MLLMs calculate loss solely based on textual tokens, limiting the flow of information from other modalities. In contrast, recent advancements in multimodal generation incorporate image data into the optimization process, enabling richer cross-modal interactions. Focusing on dynamic spatial reasoning, Li et al. [24] compare a text-only setup with an interleaved text-image scenario where images do not contribute to the loss. Interestingly, even when the model is optimized solely on text tokens, interleaved text-image training outperforms or matches text-only training. When supervision signals are derived from both text and visual tokens, the model demonstrates superior performance, generating more effective visual and verbal reasoning traces. This highlights the importance of leveraging multimodal information to enhance supervision signals, in addition to careful loss function design.
@@ 4 Evaluation Benchmark
Robust evaluation benchmarks are essential for advancing spatial reasoning capabilities in MLLMs, as they not only measure current abilities but also highlight critical gaps and guide future research directions. Several benchmarks [84, 19, 85, 21] focused on text-only scenarios, where models were evaluated purely through language descriptions. While these benchmarks provided initial insights, the emergence of MLLMs has necessitated more comprehensive evaluation frameworks that can assess spatial reasoning across different modalities and scenarios.
@@ 4.1 Static Spatial Reasoning
VSR [31] pioneered comprehensive evaluation across 66 spatial relation types, revealing significant gaps between early VLMs and human performance, and highlighting the crucial role of positional encodings. What’sUp [43] in contrast, narrowed its scope to household object pairs in spatial relationships. It uncovered models’ preference for color attributes over spatial relationships, highlighting specific limitations in spatial understanding. Similarly, CV-Bench [74] evaluated the fundamental 2D (Spatial Relationship and Object count) and 3D (Depth Order and Relative Distance) visual understanding tasks of MLLMs. Beyond 2D tasks, SpatialVLM [29] introduced 3D spatial VQA tasks incorporating both qualitative and quantitative assessments, while SpatialRGBT-Bench [30] further advanced the evaluation landscape by providing ground-truth 3D annotations across diverse environments. Collectively, these benchmarks have demonstrated that using data with corresponding spatial annotations for fine-tuning can improve the model’s performance on associated tasks.
In addition to general benchmarks, other works have focused on more specialized tasks, yielding deeper insights. For instance, Wang et al. introduced three novel VQA benchmarks (Spatial-Map, Maze-Nav, and Spatial-Grid), revealing following findings: extra visual input offers limited improvement over MLLMs’ LLM backbone, and MLLMs tend to rely less on visual information when sufficient textual context is available. TopViewRS [86] specifically examined spatial reasoning from top-view perspectives, revealing a substantial performance gap between MLLMs and humans. Additionally, COMFORT [23] systematically assessed Frames of Reference (FoR) using 3D images. Their findings showed that current MLLMs struggle with robustness and consistency in FoR tasks. Specifically, most models exhibit a preference for reflected coordinate transformations and egocentric relative FoR but face challenges in adapting flexibly to other FoRs. 3DSRBench [87] assesses MLLMs’ 3D spatial reasoning capabilities and reports that even state-of-the-art models perform poorly and scaling laws for MLLMs are not effective for 3D spatial reasoning.
@@ 4.2 Dynamic Spatial Reasoning
The recent work SAT [11] goes beyond static questions to more dynamic tasks, including Object Movement, Goal Aiming, Action Consequence, Egocentric Movement and Allocentric Perspective. The results demonstrated that stronger closed-source models (GPT4-o [1] and Gemini-1.5-pro [88]) and spatially-tuned models (Robopoint [89] and Cambrian-1 [74]) still struggle on dynamic tasks despite performing well on static. In particular, the perspective tasks involving Transformation Reasoning capability remain challenging even for strong models like GPT4-o. iVISPAR [90] introduces an interactive visual-spatial reasoning benchmark designed to systematically evaluate MLLMs as agents in dynamic environments. Experimentation results demonstrate that while state-of-the-art VLMs can handle basic spatial reasoning tasks, they face significant difficulties with more complex scenarios, especially in 3D environments.
Another recent study VSI-Bench [13] evaluated dynamic spatial reasoning tasks through a video-based benchmark. One of their key observations is that the best-performing MLLM use global spatial context and reasoning to infer correctly, indicating MLLMs may develop an implicit world model. And the error analysis reveals that almost 50% of errors stems from egocentric-allocentric transformation, indicating a lack of Transformation Reasoning capability. Additionally, linguistic prompting techniques like CoT [91], although effective in language reasoning and general visual tasks, are harmful for spatial reasoning tasks. They also leveraged cognitive maps [92] to explore how MLLMs remember spaces, finding it forms a series of local world models from the video, rather than a unified global model. In addition, SPACE [41] and VisFactor [93] evaluate whether MLLMs demonstrate spatial cognition and find that even frontier models perform near chance level.
Despite the comprehensive evaluation of MLLMs’ relational and transformation reasoning abilities across various benchmarks, there remains a gap in correlating these abilities with real-world applications, such as Embodied AI. While current benchmarks provide valuable insights, they lack empirical and theoretical investigations that demonstrate how better spatial reasoning improves application performance. We believe that future research should bridge this gap, providing a deeper understanding of how stronger spatial reasoning in MLLMs can lead to better outcomes in practical applications.
Our position:  Existing benchmarks lack a direct correlation with real-world applications like Embodied AI. We advocate for an evaluation framework that provides investigations to establish a connection between enhanced spatial reasoning in MLLMs and improved application performance.
@@ 5 Reasoning Method
Direct reasoning involves the model directly generating the answer. Instead of directly outputting the answer, recent studies have shown that scaling test-time compute [94, 95] improves reasoning performance. This is achieved through methods such as search algorithms or strategies like CoT reasoning [91, 96]. Within spatial reasoning, the effectiveness of CoT reasoning has also been extensively explored. For example, Li et al. [86] demonstrated its utility in eliciting spatial reasoning. Besides, Wu et al. enhanced performance in grid-based dynamic spatial reasoning tasks by representing reasoning traces in an ASCII-style format. SpatialCoT [98] improves action generation by introducing a two-stage framework that first aligns inputs with spatial coordinates and then performs chain-of-thought spatial grounding. However, in multimodal spatial reasoning, textual descriptions can be verbose and sometimes imprecise in describing the spatial relationships in the image. MVoT [24] introduces a novel reasoning paradigm which extends reasoning traces from textual to multimodal representations. This opens new direction and opportunities for multimodal native models to tackle the problem of spatial reasoning, with more straightforward illustrations during reasoning. Reinforcement learning has been applied across a wide range of vision-related tasks [99, 100], especially given the rise of GRPO as in DeepSeek-r1 [101]. Visual Planning [102] takes a step forward in using RL to optimize visual-centric planning trajectory, achieving better performance than language-reasoning baselines. Nonetheless, the role of RL in improving spatial reasoning remains underexplored, highlighting a valuable avenue for future investigation.
In addition to utilizing the reasoning capability of a single model, recent advancement in multimodal agentic reasoning unlocks the potential of incorporating multiple models, each acting as an expert in specific tasks, to improve spatial reasoning performance. For example, Image-of-Thought prompting [103] introduces a spatial ruler to accurately perceive spatial relationships and address spatial positioning queries. Similarly, SpatialBot [104] enable vision-language models (VLMs) to call an external API for depth estimation, enhancing the understanding of spatial information in input images. MM-ReAct [105] employs vision expert models within a ReAct-style pipeline to gather observations that support decision-making. Despite these efforts, research on designing agentic workflows specifically for spatial reasoning remains limited. Future directions could explore the agentic workflow tailored for spatial reasoning, for example, either by structuring the spatial relations with scene graph [106], or by imagining the spatial transformation with MVoT-alike systems.
@@ 6 Case Study
Figure 2: The responses and attention visualization of Qwen-2VL on path reading and planning tasks.
To further illustrate the limitations of current MLLM, we designed two representative tasks, as illustrated in the top section of Figure 2. In the Path Reading task, the model is provided with an image of a maze and asked to describe the path in it. The Planning task goes a step further by requiring the model to plan a path from a given start point to a specified goal in the maze. The detailed prompts can be found in Appendix A. In this case study, we employ Qwen2-VL [5] and LLaVa 1.5 [68]. Inspired by ADAPTVIS [107] and ViCrop [108], we additionally analyze the attention distribution over the image to gain deeper insights into the model’s spatial reasoning capabilities. The results of Qwen2-VL are shown in Figure 2 with LLaVa 1.5 [68] results in Appendix A.
The goal of the Path Reading task is to evaluate whether the model can accurately attend to the corresponding regions in the image. Although both models fail to generate a completely correct answer, our findings reveal distinct attention patterns between the two models. Qwen-2VL, compared to LLaVA-1.5, demonstrates a more accurate attention mechanism, as its correct responses consistently focus on the relevant regions of the image. This observation aligns with the findings of ADAPTVIS, which reported a strong correlation between successful spatial reasoning and the model’s ability to align its attention distribution with actual locations. The difference between these models may be attributed to architectural enhancements in Qwen-2VL, such as the use of 2D-RoPE in the vision encoder and M-RoPE in the language model backbone. These findings further support our analysis of position embedding in Sec 3.2.
The Planning task requires more complex spatial reasoning capabilities. Our results indicate that both Qwen-2VL and LLaVA-1.5 fail to generate correct path, while responses of Qwen-2VL are much more reasonable. The attention maps demonstrate that Qwen-2VL primarily focuses on object-level semantic information, such as ‘yellow star’, rather than forming a higher-level understanding of the spatial layout. To further explore these limitations, we introduced extra visual cues, as illustrated in the bottom section of Figure 2. Additionally, we employed the CoT prompting technique to investigate the reasoning steps. The results reveal that Qwen-2VL generates a reasonable path while ignoring the black boundary. As shown in attention maps, Qwen-2VL also attends to the visual cues during its reasoning process. However, it remains challenging when reasoning about the accurate spatial relationship, like ‘right to the red triangle’.
@@ 7 Conclusion
This position paper discusses a critical and timely gap of spatial reasoning capabilities in MLLMs. By systematically dissecting the challenges in spatial reasoning, we underscore the urgent need for targeted research and development. Our analysis reveals that the components of current recipes-from training data to reasoning mechanisms-limit the spatial reasoning capability of MLLMs, not only constraining real-world applicability but also hindering the way to true intelligence. Therefore, we call on the community to prioritize spatial reasoning and develop new recipes for MLLMs to effectively enhance their spatial reasoning capabilities.